D' Souza, Inc., issues $900,000 of 11% bonds that pay interest semiannually and mature in seven years. Assume that the market interest (yield) rate is 12% per year compounded semiannually. Compute the bond issuance price.
Answer #1
(note there will be slight difference in issue price calculation ,kindly use the factors as given
inn your question to get exact answer)
Bond characterstics Amount
1-a) Principal 900,000
interest              49,500
Market interest rate 6%
periods to maturity 14
issue price 858,172 answer
Calculation of bond issue price
Where
i= 6.00%
t= 14
principal * PV of $1 at 6% for 14 yrs =
900,000 * 0.4423        = 398070
interest * PV of ordinary annuity at 7%=
49500 * 9.29498 = 460102
bond issue price 858172